Suppose that 10 years ago, you purchased shares in a certain corporation's stock. Between then and now, there was a 3:1 split and a 5:1 split. If shares today are 71% cheaper than they were 10 years ago, what would be your rate of return if you sold your shares today? Round your answer to the nearest tenth of a percent.
Answer #1
For this question, we need to first calculate the no of shares now available for sale in year 10.
And also, we need to calculate the share price after year 10.
Assumed share price at T0= $      100.00
Share purchased at T0=               1.00
Initial investment $      100.00
Share price at T10= 100*(1-71%)
Share price at T10= $         29.00
Share purchased at T0=               1.00
First split= 3:1
share available after 1st split=               3.00
2nd split= 5:1
share available after 2nd split= 5*3
share available after 2nd split=             15.00
Share selling price at T10= Share for sale * Share price at T10
Share selling price at T10= 15*29
Share selling price at T10= $      435.00
Rate of return= (Share selling price - Initial investment)/Initial investment
Rate of return= (435-100)/100
Rate of return= 335.00%
The annualized rate of return= ((Share selling price at T10/Initial investment)^(1/time))-1)
The annualized rate of return= ((435/100)^(1/10))-1)
The annualized rate of return= 15.84%